A snowboarder is attempting to clear a gap after a jump, and it is known that he must  be travelling at least 34 mph to clear the gap. If the distance from the start of hill to the  jump is 42 yards long, what is the acceleration of the snowboarder if he starts from  rest? Please express your answer in ft/s2. (Hint: 1760 yards = 1 mile)


1
Expert's answer
2021-06-07T15:01:39-0400
"v^2=2as\\(34\cdot1.46667)^2=2(42\cdot3)a\\a=9.87\frac{ft}{s^2}"